OpenStudy (anonymous):

Find The Solutions Of The Equations -5x=-21+6y -5x+1y=-84

OpenStudy (anonymous):

-5x = -21 + 6y -5x = -84+1y

OpenStudy (ajprincess):

Subtract the second equation from the first equation and plz tell me vat u get @DamnImBad.

OpenStudy (ajprincess):

-5x=-21+6y -5x=-84-1y

OpenStudy (anonymous):

x=15 y=-9

OpenStudy (ajprincess):

yup perfect:)